OpenStudy (anonymous):

Can anyone help with the finding an arbitrary k value in a system of equations. x-y=k-1 2x-2y=k find the values where the system will have no solution,infinite solution, and one solution

OpenStudy (anonymous):

Multiply the 1st eqn with 2 and then subtract the the 2nd eqn from the first. The remaining eqn should be 0=2k-2-k...0=k-2...k=2